Systems Staff Meeting Notes 6/5/00

Systems Staff Meeting Notes
June 5, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Everett has accepted an internship with Disney at the Animal Kingdom from June 19 - January 4. He should be returning; we will attempt to hire a techie for 6 months in his absence.
  2. DLC Programmer/Coordinator position - received 5 resumes, 3 were not qualified - will interview the 2 candidates this week. Would like to have it all wrapped up by next week.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Another increment of barcode scanners has been ordered.
  2. 16 Dells are in the pipeline.
  3. Another order of 10 Dells has gone out.
  4. 5 Gateway laptops are order.
  5. 2 dozen Palm Visors are order from Best Buy.
  6. PS 16/600 printers - repair or replace? LAC has one with a broken paper feed. Bill says to give them a loaner off the shelf and we'll order them a new HP and JetDirect.
Liaisons
  1. Everett chaired an AD HOC liaison training committee. They submitted their report to Martha on Friday and will present it to the liaisons on Thursday.
Networking
  1. Will discovered that NERDC had misconfigured the MSL IPX feed back in February. On Friday they reset the router during midday. They had it set for the wrong frame type. It was not affecting the whole building, just certain parts.
  2. There was a Xerox IPX outage Sunday night at MSL. There was no trouble accessing the CD server, only the print server. Somehow the print release stations did not get hooked back up to the mothership. It appears to be working this morning.
  3. With the recent power outage in East a few weeks ago, Bill discovered that the generator located out back by the loading dock actually belongs to Criser. There is also another back there that is used by Matherly. We have no source of backup power for East. Bill will look into this matter.
  4. NT9 - has been problematic since installing the hot fix two weeks ago. Will noticed a NetShield conflict as well. He removed the virus software and the hot fix which seemed to resolve the printing (and most of the permission) problems. The DLC permission problems still exist. Will is going to try and put the virus scan back on, but the Soda Shoppe has changed their procedures for downloading it. He'll try to resolve this sometime today.
  5. Resource Services felt horribly inconvenienced by the power outage that affected their printing. We're looking at an alternative - adding print queues to their NT work stations. It would require 2 queues - they would have to keep up with which is which. Will look at Somba & Linux if this doesn't work.
  6. There was an attachment problem with Outlook on Friday. Scanmail blocked access to all attachments. Bill turned off the real time scan. It is still finding some viruses, would like to get it going again. We'll reinstall/de-install when we receive the full license.
  7. Gnutella virus - generated via Gnutella distribution. Gnutella is an underground (with no central server) replacement for Napster.
  8. West wiring was supposed to have been finished by Digital over the weekend.
  9. Will received the Novell CD conversion software on Friday. He'll try to install today or tomorrow.
  10. Wireless - waiting on Facilities to install the conduit so that Digital can install the wiring.
Software Projects Status
  1. New project started today - Conservation Area for Preservation. The data is used by John Freund (Special Collections). Will have to migrate the data from Access to SQL.
  2. Preservation Q&A - it is pretty much finished with some basic coding. Waiting for Nelda to return from vacation before migrating the data.
  3. Personnel System - modifications are finished.
  4. Library Statistics - changes made, it is complete. Will check with Cecilia regarding use at the Health Center.
  5. Hilton has just about completed the modifications on LASSIE. Marie Kennedy will be leaving for a year and Michelle Crump will be taking over responsibility for LASSIE.
DP Coordinator
  1. New student, Andrew Zahn, has been hired as Lindsay's replacement. He'll start on Wednesday, shadowing Lindsay for her last two weeks.
  2. Debra brought about 95 tapes back from NERDC that we had stored over there for over 450 days without activity. They have been returned to their racks in the tape vault.
  3. Debra gave Betty Mitchell a prediction of 12 jugs of water to last us until the end of July.
Calendar
  1. Everett's last day is Thursday, June 15.
  2. Winston out June 19-23.
  3. Bobbie out July 10-21.
  4. Will out July 24-28.


Back to the Systems Homepage

Last updated June 12, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu